OpenStudy (anonymous):

REALLY NEED HELP!!! :) Line AB passes through A(-3, 0) and B(-6, 5). What is the equation of the line that passes through the origin and is parallel to line AB ?

OpenStudy (anonymous):

5x – 3y = 0 -x + 3y = 0 -5x – 3y = 0 3x + 5y = 0 -3x + 5y = 0

OpenStudy (anonymous):

I really have no clue how to do this :((((

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

they are asking you for the slope of the line or at least you need that first

OpenStudy (anonymous):

Can you walk me through i'm completely lost

OpenStudy (misty1212):

you know the slope of the line between (-3, 0) and (-6, 5) ?

OpenStudy (misty1212):

if you do not, i can show you how to find it

OpenStudy (anonymous):

yes please help me find it

OpenStudy (misty1212):

first we think, then we do it the math teacher way

OpenStudy (misty1212):

from -6 to -3 is right 3 units from 5 to 0 is down 5 units right 3, down 5, slope is \(-\frac{5}{3}\)

OpenStudy (misty1212):

now the math teacher way \[\frac{y_2-y_1}{x_2-x_1}=\frac{5-0}{-6-(-3)}=\frac{5}{-3}=-\frac{5}{3}\]

OpenStudy (anonymous):

Okay so how is finding the slope going to give me the answer?

OpenStudy (misty1212):

then since it goes through the origin \((0,0)\) the point slope formula gives \[y=-\frac{5}{3}x\]

OpenStudy (misty1212):

multiply by 3 to get \[3y=-5x\]

OpenStudy (misty1212):

add \(5x\) to both sides to put it in standard form get \[5x+3y=0\]

OpenStudy (anonymous):

oooooh okay :) Thank you!!!!

OpenStudy (misty1212):

i see that is not an answer choice though

OpenStudy (misty1212):

we can fix that multiply all by -1 get \[-5x-3y=0\]

OpenStudy (anonymous):

oh wait its not :/

OpenStudy (anonymous):

ohh alright thank you again much appreciated!

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]